8:00: we start setting up the system for calibrations.

HEX X Y Z Rx Ry
  -3.02 -6.27 4.35 -324 155

Acquired new DARK: 20141126_160957

We null the coma, new Hexapod positions:

HEX X Y Z Rx Ry
  -3.02 -7.27 4.38 -225 150

We load zero SLOPE_OFFSET, REC: 150modes_ryb_TTrm_1kHz_CoG

NOTE: the following data have been acquired with the disturbance created the 13 Nov in Arcetri. These files are just a renamed version of the ones used with RR (double pass of light on M2) so they are missing a factor 2 in amplitude when used with ARGOS calibration unit (single pass of light on M2).
TN Notes
20141126_171439 CL, no dist, G=0.5,0.4,0.3
20141126_171532
20141126_173032 CL, dist=0.8'', G=0.5,0.4,0.3
20141126_173115
20141126_173303
20141126_174242 CL, dist=1.6'', G=0.5,0.4,0.3
20141126_174311
20141126_174431
20141126_174749 CL, dist=1.6'', G=0.5,0.4,0.3, Slope_offset=20141106_195056
20141126_174817
20141126_175032

We try to close the pupil loop but the system crashes because we loose the bottom right part of the mirror (as seen from the FastGUI)

We recover the system without disturb, low gain:

TN Notes
20141126_180216 CL, no dist, G=0.1,0.1,0.0
20141126_180404 CL, no dist, G=0.5,0.4,0.3
20141126_181859 CL, dist=1.6", G=0.5,0.4,0.3, BAD

We set the Pupil motors to the default position, recorded when the IM was acquired:

PUP X Y
Blue 759 509
Yellow 285 715
Red 457 665

PUPIL TARGET: 20141126_183730

TN Notes
20141126_185229 CL, no dist, G=?

New define a new subaperture set removing the bottom right edge of the red pupil that is quite vignetted when the calibration unit is in use. The Subap removal is done automatically setting the threshold value to 0.3 of the pupil median flux value.

SUBAPERTURE SET: 20141126_192200, slope computation algorithm is CoG

We rise the diodes flux and we record a new IM for the new subaperture set

CoG calibrations with subaps 20141126_192200

LEDs @ 0.4A (~8000ADU per pupil), frame rate = 650Hz

Pupil PM X PM Y PI Rx PI RY
Blue 759 509 -0.76 0.18
Yellow 285 715 -0.54 1.89
Red 457 665 -0.55 -1.07

IM Shift REC Notes
20141126_200923 3 20141126_201400 OL
20141126_205607 4 20141126_205800 CL 10 modes
20141126_210146 3 20141126_210300 CL 50 modes
20141126_210513 4 20141126_210700 CL 150 modes
20141126_210900 4 20141126_210900 CoG REC

20141126_210900 REC test

TN Notes
20141126_211831 CL, no dist, frame rate = 650Hz, G=0.4,0.5,0.5
20141126_211909

SLOPE_OFFSET: 20141126_212039, with CoG algo

TN Notes
20141126_212100 CL, no dist, frame rate = 650Hz, G=0.4,0.5,0.5
20141126_212143

13:24: we prepare new subaperture sets with TCoG48, WWPA and CoGNorm algorithms:

SUBAPERTURE SET: 20141126_212700, slope computation algorithm is TCoG48

SUBAPERTURE SET: 20141126_212800, slope computation algorithm is WWPA

SUBAPERTURE SET: 20141126_212900, slope computation algorithm is CoGNorm

TCoG48 calibrations with subaps 20141126_212700

Pupil PM X PM Y PI Rx PI RY
Blue 759 509 -0.76 0.18
Yellow 285 715 -0.54 1.89
Red 457 665 -0.55 -1.07

IM Shift REC Notes
20141126_214133 4 20141126_214200 OL
20141126_214432 4 20141126_214600 CL 10 modes
20141126_214755 4 20141126_214800 CL 50 modes
20141126_215027 4 20141126_215200 TCoG48 REC

20141126_215200 REC test

SLOPE_OFFSET: 20141126_215647, with TCoG48 algo

TN Notes
20141126_215928 CL, no dist, frame rate = 650Hz, G=0.4,0.5,0.5
20141126_220009
20141126_220059
20141126_220434 CL, dist= 1.6", frame rate = 650Hz, G=0.4,0.5,0.5, BAD
20141126_220558 CL, dist= 1.6", frame rate = 650Hz, G=0.4

Measuring on the fly the slope RMS we get the following values: 62nm rms in OL, 23 nm rms with G=0.4 to all modes, 33nm rms with G=0.1

14:50: still working with calibration unit

ADSEC FLAT: 20141126_argos150modes: acquired closing the loop on 150modes (with TT) measured by the LGSW.

PBL with BCU: loading a new slope_offset stacks the BCU. We have to restart the basdard to recover.

TN Notes
20141126_232652 CL, dist=1.6", frame rate = 650Hz, G=0.4,0.5,0.5
20141126_232738
20141126_234312 REC CoG, G=0.4,0.5,0.5
20141126_234348

15:40: we start the calibrations for the CoGNorm algorithm but we have to stop soon...

JITTER TARGET: 20141127_004700: with positions used in the TCoG48 and CoG calibrations

CoGNorm calibrations with subaps 20141126_212900

Pupil PM X PM Y PI Rx PI RY
Blue 759 509 -0.76 0.18
Yellow 285 715 -0.54 1.89
Red 457 665 -0.55 -1.07

IM Shift REC Notes
20141126_234639 4 20141126_214800 OL
20141126_235133 4 20141126_215300 CL 10 modes

Pockels Cells Transmission test

Sweep between 0.1 and 1.5us with 4 steps with Internal Sources.

The system configuration is the following:

Pupil PM X PM Y PI Rx PI RY
Blue 759 509 -0.76 0.18
Yellow 285 715 -0.54 1.89
Red 457 665 -0.55 -1.07

TN Sup
'20141127_012314', '20141127_012439' B=77, Y=100, R=98
'20141127_012646', '20141127_012939' B=74, Y=104, R=102 (7 steps)

-- MarcoBonaglia - 09 Dec 2014
Topic revision: r2 - 09 Dec 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback